# Particle Identification with MLPs and PINNs Using HADES Data Marvin Kohls1

### v1

@@ 2.2 Charged Hadron Identification with Neural Networks
with the classification (binary) cross-entropy loss , the two terms with single correct label cross-entropy for the domain classifier and the (binary) cross-entropy loss for the specific energy loss.
@@
We focus on the reconstruction of -mesons without applying preselection criteria. These mesons are rare and only produced once in every 100th event in the specific presented case of Ag(1.58 A GeV)+Ag collisions.
@@ Acknowledgements
This research was supported by GSI Helmholtzzentrum für Schwerionenforschung GmbH. The author also gratefully acknowledges the HADES Collaboration for providing the data, and Simon Spies and Waleed Esmail, whose earlier work on neural network–based analyses of HADES data formed an inspirational foundation for this study.

### v2

@@ 2.2 Charged Hadron Identification with Neural Networks
where implements the Bethe-Bloch formula for charged Kaons ( MeV/) in the MDC gas mixture (70% Ar, 30% CO2)111The physics penalty is implemented as a custom Keras layer that calculates the theoretical d/d from momentum and adds the MSE between measured and expected values with .
@@
Correspondingly, we focus on the reconstruction of -mesons without applying preselection criteria. These mesons are rare and only produced once in every 100th event in the specific presented case of Ag(1.58 A GeV)+Ag collisions.
@@ Acknowledgements
This research was supported by GSI Helmholtzzentrum für Schwerionenforschung GmbH. The author also gratefully acknowledges the HADES Collaboration for providing the data, and Simon Spies and Waleed Esmail, whose earlier work on neural network–based analyses of HADES data formed an inspirational foundation for this study. All figures are classified as original work.